PLEASE HELP AND SHOW ME HOW TO DO THIS THANK YOU SO MUCH

F Enterprises

Balance Sheet (partial)

At December 31, Year 4

Stockholders' Equity:

Preferred stock ($75 par)

$150,000

Common stock ($0.05 par)

7,500

Additional Paid in capital

2,246,500

Total paid in capital

2,404,000

Retained earnings

375,000

Treasury stock (4,000 common shares)

(72,000)

Total stockholders' equity

$2,707,000

1. How many shares of preferred stock have been issued?      ______

2a. How many shares of common stock have been issued?     ______

2b. How many of the common shares are outstanding?          ______

3. If the common shares were issued at $15 per share, at what average price per share were the preferred shares issued?             $_____

4. If retained earnings at the beginning of the period was $300,000 and $125,000 was earned in net income, what was the amount of dividends declared during the year?$______

5. What was the average cost per share of the treasury stock acquired? $_______
Answer #1

1) Preferred stock issued = 150000/75 = 2000 Shares

2a) Common Stock issued = 7500/.05 = 150000 Shares

2b) Outstanding shares = 150000-4000 = 146000 shares

3) Average issue price = (2404000-150000*15)/2000 = 77 per share

4) Dividend = 300000+125000-375000 = $50000

5) Cost per share = 72000/4000 = 18 per share